OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

answer chocies

OpenStudy (anonymous):

There aren't any :/

OpenStudy (anonymous):

are u doing algebra

OpenStudy (anonymous):

Haha yesss. Algebra 1.

OpenStudy (anonymous):

Can you help me with this? do you know itt?

OpenStudy (anonymous):

hang on i have a Algebra book

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

do u have to pick one of those

OpenStudy (anonymous):

Noo all. :/

OpenStudy (anonymous):

i past Algebra but wat are u doing

OpenStudy (anonymous):

wit n a+

OpenStudy (anonymous):

what do you mean what am i doing? im trying to solve this....

OpenStudy (anonymous):

u half to explain the soultion using complete sentences

OpenStudy (anonymous):

Ik that..

OpenStudy (anonymous):

make two equations using complete sentences

OpenStudy (anonymous):

make a math problem but put it in sentences

OpenStudy (anonymous):

thanks for trying to help. Im going to go now though.! byee

OpenStudy (anonymous):

for two lines to be intersecting, their slopes should NOT BE EQUAL they can be perpendicular! so, consider two lines in slope-intercept form \[l_1\equiv y=x+b_1\\ l_2\equiv y=-x+b_2\] let the two lines intersect at some common point (1,2) solving for b1, and b2: \[2=1+b_1\implies b_1=1\\ 2=-1+b_2\implies b_2=3\] so the two intersecting lines are : \[y=x+1\\ y=-x+3\]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

Ohhh thanks :) and whats kapeesh??

OpenStudy (anonymous):

Yes, me kapeesh hahaha.!